Circular No. 7789

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

IM NORMAE W. Liller, Viña del Mar, Chile, reports that he has detected an ap-

parent outburst of the old nova IM Nor (Woods and Bailey 1920, HCO Bull. 734) at red mag 8.3 on Technical Pan photographs taken during Jan. 10.343–10.360 UT with an 85-mm f/1.4 lens (+ orange filter); the star was subsequently found at mag ≈ 8.5 at the very edge of another pair of photographs taken on Jan. 3.348 and 3.349. Nothing was visible at the position of this star on 2001 Oct. 13 (limiting mag 11.5). A CCD image taken by Liller on Jan. 12.330 with a 0.20-m f/1.5 Schmidt camera (with a minusred-and-infrared filter) yields broadband V = 8.13 ± 0.03 and the following precise position for the nova: α = 15^{h}39^{m}26^{s}.61 ± 0^{s}.18, δ = −52^{o}19′18′′.6 ± 0′′.5 (equinox 2000.0); Woods’ position from a 1920 plate is α = 15^{h}39^{m}27^{s}, δ = −52^{o}19′.4. A low-dispersion objective prism spectrum, taken by Liller on Jan. 12.351, shows Hα to be very strong and relatively narrow with FWHM = 950 ± 50 km/s. Other lines of the Balmer series, the Fe II multiplets 42 and 74, and O I at 777.1 nm are also clearly present. Elliot and Liller (1972, Ap.J. 175, L69) suggested a possible link between the x-ray source 2U 1536–52 and IM Nor, and they found no other outburst of the nova on Harvard plates taken during 1894–1954. No known spectroscopic observations of IM Nor were previously ever made.

SUPERNOVA 2002A IN UGC 3804 S. Benetti, G. Altavilla, A. Pastorello, M. Turatto, S. Desidera, R.

Brunino, M. De Checchi, and S. Villanova, Padova; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, write: “A spectrum (range 400– 780 nm, resolution 2.5 nm) of SN 2002A (cf. IAUC 7786), obtained on Jan. 11.12 UT with the Asiago 1.82-m telescope (+ AFOSC), shows that it is a type-IIn supernova soon after explosion. The spectrum is dominated by narrow (FWHM 1200 km/s) Balmer emission lines with broader wings. The spectrum also shows a relatively blue continuum. A recession velocity of 2650 km/s is derived from the emission-line positions.”

P. Antonini, Avignon, France, reports that SN 2002A is present on an unfiltered CCD image taken with a 0.3-m reflector on Jan. 9.99 UT at mag ∼ 16.6 and position α = 7^{h}22^{m}35^{s}.46, δ = +71^{o}35′41′′.8 (equinox 2000.0)

PECULIAR VARIABLE IN MONOCEROS Corrigendum. On IAUC 7785, line 8, for Oksanan read Oksanen

2002 January 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green